OpenStudy (anonymous):

how do i do 3x + 2y = 8 and find the points it will be on a graph?

OpenStudy (anonymous):

Well one way is solve for y 3x + 2y = 8 Subtract 3x on both sides 2y = -3x + 8 Divide by 2 on both sides y = -3/2x + 8/2 y = -3/2x + 4 So slope is m in the equation y=mx+b So your slope is -3/2

OpenStudy (anonymous):

thanks